# On Joint Noise Scaling in Differentially Private Federated Learning with Multiple Local Steps

### v1

@@ 5 Experiments
In this section we empirically demonstrate the benefits of being able to run more than a single local optimisation step under limited communication budgets. Our chosen settings mimic a typical cross-silo FL setup: there are a limited number of clients, each having a smallish local database. The clients have enough local compute to run optimisation on the chosen model, while the server-client communications required for updating the global model are the main bottle-neck.
In this setting, it makes sense to try and push more optimisation steps to the clients while reducing the number of global updates (FL rounds). We also assume that the clients send their local updates via some trusted aggregator (which we only assume and do not implement in practice in the experiments). For more details on all the experiments, see Appendix Appendix.
We first train a small convolutional neural network (CNN) on Fashion MNIST data [47], that is distributed iid among 10 clients. We use the CNN architecture introduced by [35] (see also [43]). Figure 1 shows the mean with standard error of the mean (SEM) over 5 repeats for test accuracy and loss with DP-SGD using Skellam noise with 32 bits gradient quantisation,.i.e., without quantisation [1]. We train the model for 20 FL rounds and varying number of local steps. Comparing the results for a single local step as opposed to $1$ local epoch, it is evident that being able to take more local optimisation steps (as allowed by Theorem 4.7) can bring considerable utility benefits under a fixed privacy and communication budget.
@@
Figure 1: CNN on Fashion-MNIST, 10 clients, mean and SEM over 5 seeds, Skellam mechanism, unbounded $(\varepsilon=1,\delta=1e-5)$-DP.
Overall, assuming fixed communication and privacy budgets, we would expect the benefits from being able to run more local steps to be more accentuated with more complex models, while for simple-enough models even a few local steps might give good results. To test to what extent this is true for simple yet still useful models, we consider CIFAR-10 data [28]. Similar to [43], we take a ResNeXt-29 model [48] pre-trained with CIFAR-100 data [28], remove the final classifier, and use it as a feature extractor to transform the input data. We distribute the transformed CIFAR-10 data iid to 10 clients, and train linear classification layers from scratch for $10,20,40,80$ and $160$ FL rounds using DP-SGD with Skellam noise, 32 bit gradient quantisation, and varying number of local steps. Due to heavy computational costs, we use a single batch size for each configuration of local steps (concretely, we pick the best batch size value from the grid $\{64,128,256,516\}$ when using Bayesian Optimisation to tune the other hyperparameters with $20$ FL rounds. This results in choosing batch size $128$ for $1$ local step and $256$ for $1$ local epoch), and optimize all other hyperparameters separately for each FL round setting.
The mean and SEM over 5 seeds of the best results for each model over the training run are shown in Figure 2. The benefits of being able to run more than a single local steps are again clear; even with the relatively simple linear model, using $1$ local step needs roughly an order of magnitude more FL rounds to reach a similar performance compared to using $1$ local epoch.
@@
Figure 2: Mean and SEM over 5 seeds of the best performance over training runs for Linear models on CIFAR-10 using pre-trained ResNeXt29 as feature extractor for varying number of FL rounds, 10 clients, Skellam mechanism, unbounded $(\varepsilon=1,\delta=1e-5)$-DP.
To further test the robustness of the possible benefits from being able to run more than a single local optimisation step, we train even simpler Logistic Neural Network (LNN) model (i.e., 1-layer fully connected linear classification network similar to the one used in the previous experiment, but without any pre-trained feature extractor) on ACS Income data [15]. Unlike the synthetic iid data splits used in the previous experiments, Income data has an inherent client split corresponding to $51$ states from where the data has been collected. Since the inherent split is heterogeneous (different states have very number of samples as well as different data distributions), we would expect the benefits of doing more local optimisation steps between global communication rounds to dwindle, since the local models from different clients could diverge when only trained locally. However, as shown in Figure 3, even in this setting taking more local steps is very beneficial. This clearly demonstrates the utility of our analysis.
@@
Figure 3: LNN on ACS Income, 51 clients, mean and SEM over 5 seeds, Skellam mechanism, unbounded $(\varepsilon=1,\delta=1e-5)$-DP.
@@
Finally, our results might sometimes be of use also outside standard FL: consider a setting where we have $N$ copies of a model trained on disjoint data sets (e.g., one could think of independent parties learning a classifier on top of a common pre-trained model as in the previous experiment or fine-tuning a common pre-trained model; either scenario would usually lead to shared model structure and possibly also hyperparameters without explicit coordination), and the parties would like to combine the models post-hoc without running any joint training from scratch. Since this can be seen as FL with a single FL round, if the original model training on each party satisfies Assumption 4.1 (or the relaxed assumptions in Appendix Appendix), then a simple averaging of the weights will result in a joint model with improved privacy guarantees against adversaries without access to the original models.
@@
To demonstrate this effect, we account for privacy assuming the same linear model used in Figure 2 (but without actually training any models), Skellam mechanism with $32$ bit gradients, Poisson subsampling with sampling probability $0.1$, and varying number of parties and local steps. The accounting is done as it would be done in a realistic setting: we first find a noise level $\sigma_{LDP}$ that results in the target privacy level (unbounded $(\varepsilon=5,\delta=1e-5)$-LDP) for each separate model with the chosen number of local steps. We then assume that the local training satisfies Assumption 4.1 and calculate the privacy for averaging varying number of local models. As shown in Table 1, combining even $2$ models results in clearly improved privacy for the averaged model.

### v2

@@ 5 Experiments
Setup and Motivation:
Our chosen settings try to mimic a typical cross-silo FL setup: there are a limited number of clients, each having a smallish local database. The clients have enough local compute to run optimization on the chosen model, while the number of server-client communications required for updating the global model are the main bottle-neck. Note that this bottleneck will emerge even with larger actual organisations training models with broadband connections, when the model size is large-enough, e.g., when training foundation models (Bommasani et al., 2021). This is especially true when using SecAgg protocols, since the cost of running a real SecAgg algorithm presents significant compute and communication overheads even with the efficient protocols discussed in Appendix A.4. In this setting, it makes sense to try and push more optimization steps to the clients while reducing the number of global updates (FL rounds). We also assume that the clients send their local updates via some trusted aggregator (which we only assume and do not implement in practice in the experiments. However, we do use only discrete DP mechanisms compatible with standard SecAgg algorithms in all the experiments). For more details on all the experiments, see Appendix A.6. The code for reproducing the results will be released on GitHub. [^7] https://github.com/mixheikk/DPFL-with-SecAgg-paper
CNN on Fashion-MNIST:
We first train a small convolutional neural network (CNN) on Fashion MNIST data (Xiao et al., 2017), that is distributed iid among 10 clients. We use the CNN architecture introduced by Papernot et al. (2021); Tramèr & Boneh (2021). Figure 1 shows the mean with standard error of the mean (SEM) over 5 repeats for test accuracy and loss with DP-SGD using Skellam noise (Agarwal et al., 2021) with 32 bits gradient quantization,.i.e., without quantization. We train the model for 20 FL rounds and varying number of local steps. Comparing the results for $1$ local step as opposed to $1$ local epoch ($\simeq 11$ steps, but with different sampling fraction compared to baseline), it is evident that being able to take more local optimization steps (as allowed by Theorem 4.6) brings considerable utility benefits under fixed privacy and communication budgets.
@@
Figure 1: CNN on Fashion-MNIST, 10 clients, mean and SEM over 5 seeds. Running more local steps is clearly beneficial.
Linear Classifier on Transformed CIFAR-10:
Overall, assuming a fixed privacy budget, we might expect the benefits from being able to run more local steps to be more accentuated with more complex models and very limited communication budget, while for simple-enough models and more FL rounds even a few local steps could lead to good results. To test to what extent this is true for simple yet still useful models, we consider CIFAR-10 data (Krizhevsky, 2009). Similar to Tramèr & Boneh (2021), we take a ResNeXt-29 model (Xie et al., 2017) pre-trained with CIFAR-100 data (Krizhevsky, 2009), remove the final classifier, and use it as a feature extractor to transform the input data. We distribute the transformed CIFAR-10 data iid to 10 clients, and train linear classification layers from scratch for $10,20,40,80$ and $160$ FL rounds using DP-SGD with Skellam noise, 32 bit gradient quantization, and more local steps (1 epoch $\simeq 19$ steps, but with different expected batch size compared to the baseline). The mean and SEM over 5 seeds of the best results for each model over the training run are shown in Figure 2. The benefits of being able to run more than a single local steps are again clear; even with the relatively simple linear model, using $1$ local step needs roughly an order of magnitude more FL rounds over a fairly broad range of available communication budgets to reach a similar performance compared to using $1$ local epoch.
@@
Figure 2: Mean and SEM over 5 seeds of the best performance over training runs for Linear models on CIFAR-10 using pre-trained ResNeXt29 as feature extractor for varying number of FL rounds, 10 clients. Running more local steps is clearly beneficial.
Logistic Model on Income:
To further test the robustness of the possible benefits from being able to run more than a single local optimization step, we train a simple Logistic Neural Network (LNN) model (i.e., 1-layer fully connected linear classification network similar to the one used in the previous experiment, but without any pre-trained feature extractor) on ACS Income data (Ding et al., 2021). Unlike the synthetic iid data splits used in the previous experiments, Income data has an inherent client split corresponding to $51$ states from where the data has been collected. Since the inherent split is heterogeneous (different states have very different number of samples as well as different data distributions), we would expect the benefits of doing more local optimization steps between global communication rounds to dwindle, since the local models from different clients could diverge when only trained locally. However, as shown in Figure 3, even in this setting taking more local steps can be very beneficial (here $1$ epoch $\simeq 20$ steps with same local sampling fraction compared to baseline). This clearly demonstrates the utility of our analysis.
@@
Figure 3: LNN on ACS Income, 51 clients, mean and SEM over 5 seeds. Running more local steps is clearly beneficial.
Improved Privacy from Model Averaging:
@@
Finally, our results might sometimes be of use also outside standard FL. For example, consider a setting where we have $N$ copies of a model trained on disjoint data sets (for example, think of independent parties learning a classifier on top of a common pre-trained model or fine-tuning a common pre-trained model; either scenario would usually lead to shared model structure and possibly also hyperparameters without explicit coordination), and the parties would like to combine the models post-hoc without running any joint training from scratch. Since this can be seen as FL with a single FL round, if the original model training on each party satisfies Assumption 4.1 (or the relaxed assumptions in Appendix A.3), then a simple averaging of the weights will result in a joint model with improved privacy guarantees against adversaries without access to the original models.
@@
To demonstrate this effect, we account for privacy assuming the same linear model used in Figure 2 (but without actually training any models), Skellam mechanism with $32$ bit gradients, Poisson subsampling with sampling probability $0.1$, and varying number of parties and local steps. The accounting is done as it would be done in a realistic setting: we first find a noise level $\sigma_{LDP}$ that results in the target privacy level (unbounded $(\varepsilon=5,\delta=1e-5)$-LDP) for each separate model with the chosen number of local steps. We then assume that the local training satisfies Assumption 4.1 and calculate the privacy for averaging varying number of local models. Combining even $2$ models results in clearly improved privacy for the averaged model (see Table 1 in Appendix A.7).
6 Discussion
In this paper we have shown how to combine multiple local steps in DPFL using fine-grained protection granularities with SecAgg, and empirically demonstrated that this can bring considerable utility benefits under various communication-constrained settings. Our experimental results stand in stark contrasts with the message from the currently existing theoretical bounds for vanilla DPFL with FedAvg (Malekmohammadi et al., 2024, Theorem 3.2), which do not show any benefit from increasing the number of local steps. This disagreement of experimental and theoretical results underlines the need for improved theoretical analysis to understand the conditions under which increasing the number of local steps can lead to improved utility, similar to the recent breakthroughs in analysing non-DP FL (Mishchenko et al., 2022).
Acknowledgments
This research was partially supported by: 1) The Ministry of Economic Affairs and Digital Transformation of Spain and the European Union-NextGenerationEU programme for the "Recovery, Transformation and Resilience Plan" and the "Recovery and Resilience Mechanism" under agreements TSI-063000-2021-142, TSI-063000-2021-147 (6G-RIEMANN) and TSI-063000-2021-63 (MAP-6G), 2) The European Union Horizon 2020 program under grant agreements No.101096435 (CONFIDENTIAL6G) and No.101139067 (ELASTIC), and 3) the European Lighthouse on Safe and Secure AI (ELSA) from the European Union’s Horizon Europe program under grant agreement No 101070617. The views and opinions expressed are those of the authors only and do not necessarily reflect those of the European Union or other funding agencies. Neither the European Union nor the granting authority can be held responsible for them.
Appendix A Appendix
A.1 Definitions Omitted from the Main Paper
In the following, we write =$\stackrel{{\scriptstyle d}}{{=}}$ to mean equality of distributions. With this notation, we define infinite divisibility in a standard way (see, e.g., Feller 1966):
Definition A.1 (Infinite divisibility).
A distribution $F$ is infinitely divisible if for every $n$ there exists a family of distributions $F_{n}$ s.t.
$S_{n}\stackrel{{\scriptstyle d}}{{=}}\sum_{i=1}^{n}X_{i,n},$
where the random variables $S_{n}\sim F$ and $X_{i,n}\sim F_{n}$ are independent for all $i=1,\dots,n$.
A.2 Standard Algorithms Omitted from the Main Paper
For convenience, we give the pseudo-code for running DPFL with FedAvg when using DP-SGD for either sample- (Algorithm 2) or user-level (Algorithm 3) protection. We emphasize that we do not claim any novel contribution in these algorithms, but they are included simply to clarify the discussion in the main paper.
Algorithm 1 (DP)FL with FedAvg: Server-side (McMahan et al., 2017)
1:Number of clients $N$, number of FL rounds T, initial global model parameters $\theta^{(0)}$
2:for $t\in[T]$ do
3:     Choose $N_{t}$ clients for updating (w.l.o.g. numbered $1,\dots,N_{t}$) and send the current model $\theta^{(t-1)}$
4:     Receive pseudo-gradients $\Delta_{i}^{(t-1)},i=1,\dots,N_{t}$ from the updating clients
5:     Update the global model:
6:     $\theta^{(t)}\leftarrow\theta^{(t-1)}+\frac{1}{N_{t}}\sum_{i=1}^{N_{t}}\Delta_{i}^{(t-1)}$
7:end for
Output: $\theta^{T}$
Algorithm 2 DPFL: Client-side with sample-level DP-SGD using Gaussian noise (Song et al., 2013; Abadi et al., 2016)
1:Number of local optimization steps $S$, local dataset $x_{i}$, loss function $f$, learning rate $\gamma$, clipping constant $C$, noise scale $\sigma$, Poisson subsampling probability $q$
2:Client receives current model $\theta$ from the server:
3:$\theta_{0}\leftarrow\theta$
4:for $s\in[S]$ do
5:     Update local model with DP-SGD:
6:     $g_{i,s}\leftarrow 0$
7:     Sample minibatch $b_{s}$ from $x_{i}$ using Poisson sampling with probability $q$
8:     for $j\in[|b_{s}|]$ do
9:         Calculate $j$th per-sample gradient in minibatch, clip and accumulate:
10:         $\tilde{g}_{i,s}\leftarrow\nabla f(b_{s,j};\theta_{s-1})$
11:         $g_{i,s}\leftarrow g_{i,s}+\min\left(1,\frac{C}{\|\tilde{g}_{i,s}\|_{2}}\right)\tilde{g}_{i,s}$
12:     end for
13:     Add noise to clipped gradients and take a step:
14:     $\theta^{(s)}\leftarrow\theta^{(s-1)}-\gamma\left(g_{i,s}+C\xi\right)$, where $\xi\sim{\mathcal{N}}(0,\sigma^{2}I_{d})$ and $d$ is model dimensionality
15:end for
16:$\Delta\leftarrow\theta^{S}-\theta_{0}$
Output: $\Delta$ (Pseudo-gradients sent to the server)
Algorithm 3 DPFL: Client-side with user-level DP using Gaussian noise (McMahan et al., 2018)
1:Number of local optimization steps $S$, local dataset $x_{i}$, learning rate $\gamma$, clipping constant $C$, noise scale $\sigma$, Poisson subsampling probability $q$
2:Client receives current model $\theta$ from the server:
3:$\theta_{0}\leftarrow\theta$
4:Update local model $S$ times with standard SGD:
5:$\theta^{(S)}\leftarrow$ from running SGD locally for $S$ steps with learning rate $\gamma$ on $x_{i}$
6:$\Delta\leftarrow\theta^{S}-\theta_{0}$
7:Clip pseudo-gradients and add noise:
8:$\Delta\leftarrow\min\left(1,\frac{C}{\|\Delta\|_{2}}\right)\Delta+C\xi$, where $\xi\sim{\mathcal{N}}\left(0,\sigma^{2}I_{d}\right)$ and $d$ is model dimensionality
Output: $\Delta$ (Pseudo-gradients sent to the server)
Essentially, for this example when aiming for DDP, at FL round $t$ our analysis enables choosing the number of local steps $S$ in Algorithm 2 freely while still calibrating the DP noise by looking at the total noise level $\sum_{i=1}^{N_{t}}\sigma_{i}^{2}$ added by each client separately on each local step (line 14 in Algorithm 2). Note that for the special case of using Gaussian noise, a similar result has been presented by Noble et al. (2022).
A.3 Loosening Assumptions
In the main paper, as stated in Assumption 4.1, for each FL round we have assumed constant learning rate $\gamma$, norm clipping bound $C$, noise level $\sigma$, and number of local optimization steps $S$, all of them shared by all the clients. Next, we consider loosening these assumptions. As before, w.l.o.g. we consider only a single FL round, and will therefore omit the index $t$.
which can again be seen as post-processing the vector $\left(\sum_{i=1}^{N}{\mathcal{A}}_{i}^{(1)}(x_{i}),\dots,\sum_{i=1}^{N}{\mathcal{A}}_{i}^{(S)}(x_{i})\right)$, so we can again use Lemma 4.5 for accounting without encountering problems.
When considering client-specific learning rates things can be more complicated. The main issue now is to find proper sum-dominating mechanisms that satisfy:
$\sup_{x\simeq x^{\prime}}H_{\alpha}\left(\sum_{i=1}^{N}\gamma_{i,s}{\mathcal{A}}_{i}^{(s)}(x_{i})\|\sum_{i=1}^{N}\gamma_{i,s}{\mathcal{A}}_{i}^{(s)}(x^{\prime}_{i})\right)\leq\sup_{x\simeq x^{\prime}}H_{\alpha}\left({\mathcal{A}}^{(s)}(x)\|{\mathcal{A}}^{(s)}(x^{\prime})\right),s=1,\dots,S.$ (15)
As a concrete example, assume ${\mathcal{A}}_{i}^{(s)}$ is the continuous Gaussian mechanism with shared norm clipping constants and noise levels $C_{i,s}=C_{s},\sigma_{i,s}=\sigma_{s}\ \forall i$. Dropping the step index $s$ for readability, let $\gamma_{i}=\frac{\gamma_{1}}{l_{i}}$ for some $l_{i}>0,i=2,\dots,N$. Writing $g_{i}$ for a sum over the per-unit clipped gradients of client $i$, and $\xi_{i}\sim\mathcal{N}(0,C^{2}\sigma^{2}\cdot I_{d})$ a single optimization step now contributes the following term for the global update:
$-\sum_{i=1}^{N}\gamma_{i}\left(g_{i}+\xi_{i}\right)$ (16)
$=$ $-\gamma_{1}\left(g_{1}+\sum_{i=2}^{N}\frac{g_{i}}{l_{i}}+\xi\right),$ (18)
where $\xi\sim\mathcal{N}(0,C^{2}\sigma^{2}[1+\sum_{i=2}^{N}\frac{1}{l_{i}^{2}}]\cdot I_{d})$, which is a sum-dominating Gaussian mechanism. When accounting for the sum-dominating mechanism, it has sensitivity $C^{*}=\max\{C,\frac{C}{l_{2}},\dots,\frac{C}{l_{N}}\}$, which in turn gives noise variance $(\frac{C}{C^{*}})^{2}\sigma^{2}[1+\sum_{i=2}^{N}\frac{1}{l_{i}^{2}}]$ for DP.
Similarly, we could relax the assumptions further to allow the clients to use different clipping and noise levels $C_{i},\sigma_{i}$. As before, a single optimization step can again be written in the form of Equation 18, when
For global privacy accounting with a sum-dominating Gaussian mechanism, suitable sensitivity is now given by $C^{*}=\max\{C_{1},\frac{C_{2}}{l_{2}},\dots,\frac{C_{N}}{l_{N}}\}$, and the resulting variance for accounting is $\sum_{i=1}^{N}(\frac{C_{i}\sigma_{i}}{C^{*}})^{2}$.
As a simple example, assume we have 2 clients running DP-SGD: client 1 runs $S$ local steps using norm clipping constant $C$ and Gaussian mechanism with noise variance $\sigma^{2}$, while client 2 runs $2S$ local steps with clipping $C/2$ and Gaussian noise variance $\sigma^{2}$. The difference now is that while the clipping is done on each step, from the privacy accounting perspective we can disregard some noise and think that client 2 adds noise only on every other step. Looking at the update from client 2, we would then have
$=-\gamma\sum_{s=1}^{S}(g^{\prime}_{2,s}+\xi^{\prime}_{2,s}),$ (21)
where $g_{2,s}$ are the clipped per-sample gradients, $g^{\prime}_{2,s}:=g_{2,2s-1}+g_{2,2s}$, $\xi_{2,s}$ are the noise values, $\xi^{\prime}_{2,s}:=\xi_{2,2s}$, and $\mathbb{I}$ is the indicator function. Due to the clipping, the sensitivity of each fused step can be easily upper bounded via triangle-inequality: $\|g_{2,s^{\prime}}\|_{2}=\|g_{2,2s^{\prime}-1}+g_{2,2s^{\prime}}\|_{2}\leq\|g_{2,2s^{\prime}-1}\|_{2}+\|g_{2,2s^{\prime}}\|_{2}\leq C$. Since Equation 21 now has the same number of local steps as client 1 is taking, we can readily use the previous results to enable privacy accounting for the aggregated update. Combining the fusing of local steps with the previous notes on differing clipping norm values, learning rates and noise variances allows us to use our main results in several settings beyond what is stated in Assumption 4.1.
As a final note, when the clients use data subsampling for the local optimization, differing local subsampling probabilities can lead to having varying DP guarantees between the clients on the global level due to the different subsampling amplification effects, but can otherwise be incorporated with the same analysis we have already presented.
A.4 From Ideal Trusted Aggregators to Practical SecAgg Protocols
For implementing the trusted aggregator assumed in Theorem 4.6 in practice, it should be noted that as the sum over $s$ is done locally by each client during local optimization, it is always trusted as long as the individual clients are, while the sum over $i$ would need to be implemented, e.g., using a suitable SecAgg protocol. Several such algorithms are known, including the ones proposed by Bell et al. (2020); Bonawitz et al. (2017); Sabater et al. (2022); So et al. (2021).
Using a SecAgg protocol will typically also place some extra requirements on the DP mechanisms ${\mathcal{A}}_{i}^{(s)}$, since the SecAgg algorithms usually run on elements of finite rings. This precludes continuous noise mechanisms. A viable alternative is to use some suitable discrete noise mechanism, such as Skellam (Agarwal et al., 2021) or Poisson-binomial (Chen et al., 2022b). However, differing from the cases considered in the cited papers, since in our case the clients send model updates instead of single gradients, the finite ring size used in the SecAgg protocol needs to accommodate the model update size: it does no good to use Skellam mechanism with gradient quantization to a small number of bits, if the model weights and the resulting model update $\Delta_{i}$ for client $i$ still uses 32 bit floats.
A.5 Privacy Accounting Details
For privacy accounting we utilize Rényi DP (RDP):
Definition A.2.
(Mironov, 2017) Let $\alpha>1$ and $\varepsilon>0$. A randomised algorithm ${\mathcal{A}}:\,\mathcal{X}^{*}\rightarrow\mathcal{O}$ is $(\alpha,\varepsilon)$-RDP if for every $x,x^{\prime}\in\mathcal{X}^{*}:x\simeq x^{\prime}$
$D_{\alpha}({\mathcal{A}}(x)\|{\mathcal{A}}(x^{\prime}))\leq\varepsilon,$
where $D_{\alpha}$ is the Rényi divergence of order $\alpha$:
We do privacy accounting for all the experiments based on RDP. Generally, we account for the privacy of each individual local optimization step using the noise contributions from all the clients selected for a given FL round. When the clients use Poisson subsampling to sample minibatches (we assume each client uses the same probability for including any individual sample in the minibatch), we use standard RDP privacy amplification results. In practice, we use the RDP accountant implemented in Opacus (Yousefpour et al., 2021), as well as bounds for Skellam mechanism by Agarwal et al. (2021) and tight RDP amplification by Poisson subsampling (Steinke, 2022). We calculate the privacy cost of the entire training run in RDP, and then convert into ADP using (Mironov, 2017, Proposition 3). Note that, as is common in DP research, we do not include the privacy cost of hyperparameter tuning in the reported privacy budgets (see, e.g., Tramèr & Boneh 2021 for some reasoning on this practice).
A.6 Experimental Details
All the experimental settings we use satisfy Assumption 4.1. We use DP-SGD with Skellam mechanism to optimise the local model parameters, and standard federated averaging as the aggregation rule for updating the global model in all experiments. For each centralised data set (combining original train and test sets), we split the data randomly into equal shares, which results in having almost the same data distribution on each client. For hyperparameter optimization with each dataset, we first split each clients’ data internally into train and test parts with fractions (.8-.2). For tuning all hyperparameters, we use only the training fraction, and divide it further (.7-.3) into hyperparameter train-validation. We use Bayesian optimization-based approach implemented in Weights and Biases (Biewald, 2020) for hyperparameter tuning, and simulate FL using Flower (Beutel et al., 2020).
In general, when tuning hyperparameters we do 50 hyperparameter tuning runs. For each tuning run, we train the model on hyperparameter training fraction, test on the validation fraction, and try to optimise for the final model weighted validation loss. After finishing the hyperparameter tuning, we re-train the model from scratch 5 times with different random seeds with the best found hyperparameters using the entire original training data and testing on the test fraction. We report the mean and the standard error of the mean (SEM) in all the figures. In Figure 2 we plot the minimum test loss/maximum test accuracy taken over the entire training run.
For the experiments with Fashion-MNIST (Xiao et al., 2017) and CIFAR-10 (Krizhevsky, 2009) data sets, we run hyperparameter tuning separately for each combination of number of local steps {1 step, 1 epoch}, and expected minibatch sizes on the grid $\{64,128,256,512\}$ using Poisson subsampling.
For Fashion-MNIST the best expected batch sizes found are $512$ for 1 local epoch, and $128$ for 1 local step.
With CIFAR-10, due to heavy computational cost of hyperparameter tuning, we use a single expected batch size for each configuration of local steps {1 step, 1 epoch} and FL rounds $\{10,20,40,160\}$. Concretely, we pick the best expected batch size value from the above grid when using Bayesian optimization to tune all hyperparameters with $20$ FL rounds. This results in choosing expected batch size $128$ for $1$ local step and $256$ for $1$ local epoch. We then use these values and optimize all other hyperparameters separately for all other FL round settings.
With ACS Income data (Ding et al., 2021), we tune all hyperparameters for each combination of local steps {1 step, 1 epoch} with Poisson subsampling using local sampling probability on the grid $\{0.4,0.2,0.1,0.05\}$ for $10$ FL rounds. We report results on the best found local sampling probabilities ($0.05$ for both).
For ResNeXt-29 8x64, we used pre-trained weights available from https://github.com/bearpaw/pytorch-classification. Our implementation of the Skellam mechanism is based on the implementation from https://github.com/facebookresearch/dp_compression (Chaudhuri et al., 2022; Guo et al., 2023).
For American Community Survey (ACS) Income data set Ding et al. (2021) we use the data for all the states and Puerto Rico for 2018. The goal is to predict whether an individual has income greater than \$$50000$. Instead of simulating data splits, we use the inherent splits, i.e., we take each original region (state or Puerto Rico) to be a client.
For training all models, we use a small cluster with NVIDIA Titan Xp, and NVIDIA Titan V GPUs. The total compute time of all the training runs (including debugging) over all GPUs amounts roughly to 30-60 GPU days.
A.7 Additional Results
Table 1 shows the results from averaging several independently trained LDP models as described in Section 5.